Owl Shift summary 7/28/99

Mary_Manning (mem8f@virginia.edu)
Wed, 28 Jul 1999 07:06:36 -0500

Shift Summary for owl shift, 7/28/99

On Shift: Leo B, Lisa Manning

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.5
Hours of E832 data (special runs not included): 7.20
Hours of special runs (laser scans, muon runs...):0

Accounting of lost time: Tape change, Kfrend

========================================================================

Runs ended this shift: 14044

Tapes brought to FCC this shift: QK*179-181, QKB 061

========================================================================

Accesses during shift:none

========================================================================

Overview of shift: Smooth e832 running.

========================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------

14044 0:00 Take over from evening shift

394 0:55 Call MCR- SEM down to 3.9e12
SEM stays at 3.9e12 for most of run

459 2:25 Couple of HCC hot bits (1345,1617)
- experts notified

3:00 Leo notices frowny face for drift chamber plane
15. Efficiency is similar to earlier runs, but
this is the first time it is below tolerance.

3:20 Helium down to 400 psi - Dave N. notified

525 3:46 Disks full. Stop run 14044.

14045 4:04 Start 832_beam run

4:45 Leo notices small problem with etot monitoring
code which causes a frowny face

6:08 DC plane 15 is back within tolerance now.

104 6:24 Laser RMS wide - channel 2240.
Disappears next spill.

7:00 Hand over to day shift

========================================================================

Follow up Notes on Previous Shift:
No muon run, no polarity flip

========================================================================

Items to Follow up on Next Shift:

Investigate DC plane 15, HCC bits for channels 1345 and 1617.